please help me solve 4/5x=5/4
i know i need to flip the 4/5x but i keep getting confused i think the answer is
1.5 if it is im just confused on the check

Answer by stanbon(57203)

solve 4/5x=5/4
-------------
Multiply both sides by (5/4) to get:
x = (5/4)(5/4)
x = 25/16
x = 1.5625
============
Cheers,
Stan H.